Circular No. 6713

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 P1 (SOHO) C. St. Cyr, Naval Research Laboratory, reports, for the SOHO-LASCO

Consortium (see IAUC 6685), the discovery by S. Stezelberger of a another comet in SOHO-LASCO C3 coronagraph images. This comet had a tail that was > 3R_{⊙} long, and a preliminary estimate indicates that it was brighter than m_{1} = 4. Measurements by D. Biesecker and St. Cyr been reduced by G. V. Williams and are given on MPEC 1997-P04 in detail, the approximate discovery position being:

1997 UT  α_{2000}  δ_{2000}
Aug. 3.865  8^{h}26.^{m}5  +14^{o}06′

C3 observations extend to Aug. 5.700 UT, while C2 observations are from Aug. 5.634 to 5.762, at which time the solar distance was 2.7R_{⊙}. The comet is apparently a Kreutz sungrazer—in fact, second in brightness to C/1996 Y1 among the Kreutz sungrazers observed by SOHO—but it is not possible to obtain a meaningful orbit solution from the complete data set.

COMET C/1996 M1 (SOHO) St. Cyr also reports the discovery of another evident Kreutz sungrazer,

no brighter than m_{1} = 8, by Biesecker and B. McCarty in an automated search of archival SOHO-LASCO C3 coronagraph data. There was no evidence of a tail, and the comet was not found in C2 data.

1996 UT  α_{2000}  δ_{2000}
June 17.775  5^{h}41.^{m}1  +19^{o}55′

Measurements as above are given on MPEC 1997-P05, together with an orbit from all the available data.

NOVA IN THE SMALL MAGELLANIC CLOUD 1996 D. L. Welch, McMaster University, Hamilton, Ontario, reports: “The

MACHO collaboration (cf. IAUC 6312) has discovered an apparent nova in the Small Magellanic Cloud. Located at α = 0^{h}54^{m}43^{s}.04, δ = −72^{o}32′20′′.8 (equinox 2000.0), the nova peaked in brightness around 1996 June 8. However, our magnitudes around peak brightness are unreliable due to saturation. The last point before eruption was taken on May 29.77 UT. The next observation, on June 4.75, indicates that the nova had brightened by 5.4 mag. Two slope changes in the declining portion of the lightcurve occurred around June 20 and Aug. 17.”

1997 August 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden